Mccoolerson:

Maya has 5 over 6 cup of ice cream. How many 1 over 4 -cup servings are in 5 over 6 cup of ice cream? (5 points) 1 and 3 over 2 2 and 1 over 3 3 and 1 over 3 4 and 1 over 2

Mccoolerson:

thanks

Mccoolerson:

The figure below shows the quotient of fraction 5 over 8 Division sign fraction 3 over 8 using a rectangle model. Rectangle divided into eight equal parts indicating five-eighths and three-eighths, where the first three parts are shaded dark, the next two parts are shaded light, and the last three parts are not shaded. What is the quotient? (5 points) fraction 1 and 2 over 8 fraction 1 and 2 over 5 fraction 1 and 3 over 5 fraction 1 and 2 over 3

supie:

Just better way of seeing it: The figure below shows the quotient of fraction 5 over 8 Division sign fraction 3 over 8 using a rectangle model. Re-writing ^ fraction 5 over 8 Division sign fraction 3 over 8 \(\frac{5}{8}รท\frac{3}{8}\) I believe the choices would look like this because `and` means `+` or `adding` Choice 1: \(\frac{1+2}{8}\) Choice 2: \(\frac{1+2}{5}\) Choice 3: \(\frac{3+1}{3}\) Choice 4: \(\frac{4+1}{2}\) ? Also, information that was forgotten to be included?